OpenStudy (anonymous):

A rectangle is inscribed with its base on the x-axis and its upper corners on the parabola y=7-x^2 . What are the dimensions of such a rectangle with the greatest possible area?

OpenStudy (anonymous):

if the base is on the x-axis, the width of the rectangle is = x if the height is based on the parabola, the length = 7-x^2 the area of a rectangle = length * width thus, area = x*(7-x^2) = 7x-x^3 in order to maximize the area, you would need to take the derivative of the area and set it equal to 0 Area = 7x-x^3 Area' = 7 - 3x^2 7-3x^2 = 0 thus, x = 1.5275 this x represents the x needed to create the largest possible area with the given parameters. Thus: Width (x-axis) = 1.5275 Length (y-axis) = 7 - (1.5275)^2 = 4.6667 hope this helps

OpenStudy (anonymous):

ohhh okay i see thanks for the step by step

OpenStudy (anonymous):

no worries :)